OpenStudy (anonymous):

how would i write 8z^3 + 1 = 0, in the form z = re^iθ

OpenStudy (anonymous):

z^3 = -1/8 =1/8 x e^(i(pi)) z = 1/2 x e^(i(pi)/3)

OpenStudy (anonymous):

get it??

OpenStudy (anonymous):

him your answer is complex. z^3=-1/8 should have solution z=(-1/2) z=1/2*e^(i*Pi) i think

OpenStudy (anonymous):

wont you tak the cube root of the e^(i(pi)) part as well friend?

OpenStudy (anonymous):

woops both are correct

OpenStudy (anonymous):

yeah..though i can find a third root...irritating

OpenStudy (anonymous):

i figured it out, there's 3 solns : 1/2e^i(pi) ; 1/2e^i(pi/3) ; 1/2e^i(-pi/3)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks guys (:

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

its z^3=1/8*e^(i*Pi+2Pi*n) where n is any integer. we forgot the 2*Pi*n thenz= 1/2*e^((i*pi+2Pi*n)/3) for any integer n

OpenStudy (anonymous):

right